3. Suppose X ~ Beta(a, β) with the constants α, β > 0, Define Y- 1-X. Find the pdf of Y
Answer #1

X = Beta(alpha, eta )

f1-X(z) = fx(1-2)

Y = 1- X follow Beta(eta,alpha)